OpenStudy (anonymous):

3. What is the 52nd term in the sequence? 9, 12.5, 16, 19.5, 23, . . .

OpenStudy (solomonzelman):

you are adding the same number every time, correct? can you tell me what is it?

OpenStudy (anonymous):

The constant number is 3.5 @SolomonZelman

OpenStudy (solomonzelman):

yes, it is 3.5.

OpenStudy (solomonzelman):

So you can state: \(\large\color{black}{ d=3.5 }\) Now, you can see that \(\large\color{black}{ a_1}\) (the first term) is \(\large\color{black}{ 9 }\) .

OpenStudy (anonymous):

a52 = 9 + 3.5(52), right?

OpenStudy (anonymous):

Thank you very much. So now I just evaluate the equation and my answer will be given. I was just slightly confused on a1 and d in the equation.

OpenStudy (solomonzelman):

oh you were a little off.

OpenStudy (solomonzelman):

very smart, and what is \(\Large\color{black}{ a_{_{52}}=9+3.5(52-1) }\)

OpenStudy (solomonzelman):

because the formula is: \(\Large\color{black}{ a_{_{52}}=a_1+9d(n-1) }\) so you are plugging in accordingly: \(\Large\color{black}{ a_{_{52}}=9+3.5(52-1) }\)

OpenStudy (solomonzelman):

formula should have just d, not 9d, sorry.

OpenStudy (anonymous):

Can you explain briefly why the -1 is added? Is that like that in every ilteral term (or w/e its called) lol

OpenStudy (anonymous):

By the way, my answer is 187.5

OpenStudy (solomonzelman):

vconnection snapped before I could reply. Yes the answer is right...

OpenStudy (anonymous):

It's okay, and thank you. :)

OpenStudy (solomonzelman):

I'll explain this though, too. You know that (in general) \(\large\color{black}{ a_2=a_1+d }\) , yes?

OpenStudy (anonymous):

One last question, it's about common ratio's of a geometric sequence. And yes, I do.

OpenStudy (solomonzelman):

Okay, so you wanted to explain why I subtract 1, inside the formula, right? I can explain it like this: you know that: \(\large\color{black}{ a_2=a_1+d }\) \(\large\color{black}{ a_3=a_1+2d }\) \(\large\color{black}{ a_4=a_1+3d }\) \(\large\color{black}{ a_5=a_1+4d }\)

OpenStudy (solomonzelman):

so this is where we get the formula, that: \(\large\color{black}{ a_n=a_1+4(n-1) }\) right?

OpenStudy (anonymous):

oooooohhhhh, ok, I get it now somewhat. thank you again for clearing it up for me!

OpenStudy (solomonzelman):

I mean a formula that: \(\large\color{black}{ a_n=a_1+d(n-1) }\)

OpenStudy (solomonzelman):

you are adding the difference one time less than the number of the term you want to find,

OpenStudy (solomonzelman):

okay, and what would you like to know about geometric sequence?

OpenStudy (anonymous):

Here's a picture of the question: http://i.gyazo.com/a82b7d26eee146ab9dccb416834d5c36.png

OpenStudy (anonymous):

How about do you find the "common ratio" of a given geometric sequence, do you know?

OpenStudy (solomonzelman):

okay, a common ratio is a number by which you multiply (each time) to obtain the next term in a sequence.

OpenStudy (solomonzelman):

For instance, \(\large\color{black}{ a_{n-1} \times r=a_n }\) for any geometric sequence.

OpenStudy (solomonzelman):

can you tell me what pattern do you notice in the terms (of your sequence) ?

OpenStudy (anonymous):

From first glance, it looks hard. I can't really interpret the pattern.

OpenStudy (anonymous):

I do see that it's getting smaller "...-3, -18...."

OpenStudy (solomonzelman):

there is a formula: \(\large\color{black}{r=a_{n} \div a_{n-1} }\) (where r is the common ratio) basically divide any term by the term before, to find it: For example: \(\large\color{black}{r=a_{2} \div a_{1} }\) Which in your case is: \(\large\color{black}{r=(-\frac{1}{2}) \div (-\frac{1}{12}) }\)

OpenStudy (solomonzelman):

I don't see the letters working, are they working for you?

OpenStudy (anonymous):

Mhm, I see them working. Don't worry about it.

OpenStudy (anonymous):

So I can divide any two terms to get the common ratio?

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

Do they have to be in a specific order or by each other? Say if I divided -1/12 by -18, there'd be no penalty/nothing wrong with that? Or should I divide -3 by -18 (vice versa).

OpenStudy (solomonzelman):

no that would be wrong. you are dividing a term, by another term that is (right) before.

OpenStudy (solomonzelman):

Like: \(\large\color{black}{r= (a_{3}) \div (a_2) }\) \(\large\color{black}{r= (a_{4}) \div (a_3) }\) or any, \(\large\color{black}{r= (a_{n}) \div (a_{n-1}) }\) but not: \(\large\color{black}{r= (a_{3}) \div (a_1) }\) (for example.)

OpenStudy (anonymous):

Okay, so -18 by -3 would be the correct way of dividing & finding the common ratio of the sequence?

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

and what do you get the common ratio to be?

OpenStudy (anonymous):

I seeeeee now, I see. Thank you again. I'll try dividing it right now, one second. I'll tag you when I'm finished.

OpenStudy (solomonzelman):

and you can see how this actually works every time. how: ~ (-1/12) * 6 = (-1/2) ~ (-1/2) * 6 = (-3) ~ (-3) * 6 = (-18) and etc.

OpenStudy (anonymous):

r=(-1/12*-18) divided by (-1/12*3-1) @SolomonZelman

OpenStudy (anonymous):

I'm completely off lol. My answer was -5/4 (-1.25)

OpenStudy (solomonzelman):

no no R can be equal to any of the following, because all have the same value: r = (-1/12) * 6 = (-1/2) (-1/2) * 6 = (-3) (-3) * 6 = (-18)

OpenStudy (solomonzelman):

you see the 6 in there between any 2 terms?

OpenStudy (solomonzelman):

(-1/12) * 6 = (-1/2) (-1/2) * 6 = (-3) (-3) * 6 = (-18) means that r=6.

OpenStudy (anonymous):

Oh, wow. I'm looking over it now. Thank you very much @SolomonZelman I'm pretty sure I can finish the rest of it on my own now, thanks to you lol. I gave you a medal & fanned you.

OpenStudy (solomonzelman):

No need for a medal and fan, although you certainly can, if you want:) you have been a joy to teach, no kidding. You grasped it quickly! ty